I need help with math. If I have $10,000, I invest and loose 50% the first year. The second year I make 90%. What is my growth percentage? How much $ do I have?
Answer by jim_thompson5910(35256)

You lost 50% of 10,000, so you now have

10000 - 0.5*10000 = 10000 - 5000 = 5000


So after the first year, you'll have $5,000

Then you make 90%. This means that you're adding 90% of 5 thousand to 5 thousand, so...

5000 + 0.9*5000 = 5000 + 4500 = 9500


This means that after the second year, you'll have $9,500

-------------------------------------------------------

Your growth percentage is then

Percent Growth = [(New - Old)/Old] * 100

Percent Growth = [(9500-10000)/10000] * 100

Percent Growth = [-500/10000] * 100

Percent Growth = -0.05 * 100

Percent Growth = -5%

This means that you really lost 5% (and didn't grow at all)